# Guaranteeing Semantic and Performance Determinism in Flexible GPU Sharing

### 16 Mar 2026 v1

@@ 3.3 Challenge-#3: Transparency and Practicality
However, many existing spatial sharing designs [gslice, strati2024orion] sacrifice this practicality for theoretical utilization gains. Approaches relying on static compiler analysis, operator fusion, or manual kernel rewriting introduce a high "Engineering Tax". By requiring deep introspection into model graphs or invasive source code modifications, these methods tightly couple the sharing substrate with specific model implementations. This coupling creates a brittle ecosystem where even minor updates to framework or model architecture can invalidate the optimization logic, necessitating expensive manual retuning.
@@
Crucially, this coupling imposes a burden beyond development—it fundamentally expands the verification surface. In hyperscale production environments, maintaining offline-to-online equivalence is non-negotiable. Mechanisms that modify user GPU kernels or the runtime environment effectively invalidate rigorous offline validation processes. This operational reality is corroborated by infrastructure practitioners at ByteDance [bytedance_private], who note that "mechanisms requiring code rewriting are typically rejected not just for complexity, but because they mandate a prohibitive re-verification process for every framework or model update."
@@ 4 DetShare Design
This section details the design of DetShare, a system that reconciles execution determinism with high hardware utilization. Existing spatial sharing approaches often compromise semantic determinism by transparently altering block size to fit fragmented resources [coppock2025lithos]. In contrast, DetShare adopts a transparent spatial multiplexing approach. The core design philosophy of DetShare is the strict decoupling of logical execution flows from physical resource provisioning. By maintaining kernel launch configurations as immutable invariants, DetShare ensures that resource elasticity never alters the arithmetic behavior of the application.
@@ 6.1 End-to-End Performance Evaluation
We compare DetShare against industry-standard hardware partitioning mechanisms (NVIDIA MPS and MIG), as well as state-of-the-art software-defined schedulers including Orion [strati2024orion], Tick-Tock [wang2021wavelet], and Salus [yu2020salus]. Since the source code for LithOS [coppock2025lithos] is not publicly available, we implemented a faithful reproduction of its scheduling logic based on its design principles to serve as a comparative baseline.
@@ 6.1.1 Colocated AI Model Training
High Compute Contention A ResNet-50 [he2016deep]
@@ B ResNet-101
C BERT-Large [devlin2019bert]
@@ D BERT-Large ResNet-50
Low Compute Contention 1 RNN [rumelhart1986learning]
2 DeepFM [guo2017deepfm] ResNet-101
LLM Training E GLM-4-9B [zhipu2024glm] (LoRA [hu2022lora])
@@ 2 DeepFM [guo2017deepfm] ResNet-101
CV + LLM Inference A ResNet Llama 3 [llama3]
B ResNet GPT-J [wang2021gptj]
@@ 6.2.1 Traces and Baselines
To demonstrate DetShare’s transparency and effectiveness at the infrastructure layer, we evaluate an unmodified instance of the state-of-the-art serving system, SGLang [sglang2024], running entirely on top of DetShare to serve the Llama-3.1-8B-Instruct model. We compare this against a baseline of SGLang executing on a default, exclusively provisioned GPU.
@@
  • Azure LLM Trace 2024 [dynamollm2025]: Highly stochastic request arrival times, testing general adaptability.
  • LongBench [longbench2024]: Long-context tasks imposing significant pressure on GPU memory capacity and bandwidth.
  • BurstGPT [burstgpt2025]: Extreme traffic spikes challenging the scheduler’s queue management.
@@ 7 Related Work
Temporal sharing remains the default mechanism for multi-tenant GPU execution. Prior systems have significantly enhanced this approach for DNN workloads by mitigating context-switching overheads through state reuse (Salus [yu2020salus]), dynamically managing containerized or runtime memory footprints (Antman [xiao2020antman], TGS [wu2023transparent], Zico [lim2021zico]), and employing predictive scheduling to bound inference tail latency (Clockwork [clock], Gandiva [xiao2018gandiva]). However, these solutions inherently operate at the granularity of the entire GPU. Consequently, they enforce strict execution serialization and fail to overlap computation. This limitation prevents them from reclaiming idle capacity during memory-bound phases or bursty workloads, leaving significant hardware resources underutilized.
@@
Spatial sharing partitions hardware to enable concurrent execution. Foundational mechanisms include intra-context CUDA Streams [cuda:stream] and NVIDIA’s static partitioning tools, MPS and MIG. Systems such as GSlice [gslice], GPUlets [choi2022serving], and COLTI [mobin2023colti] build upon these primitives to construct higher-level spatio-temporal abstractions. Yet, because their underlying substrates rely on rigid, static partitioning, they cannot elastically adapt to the highly dynamic resource demands characteristic of modern workloads.
Recent literature pushes for even finer-grained multiplexing. POD-Attention [kamath2025pod] rewrites attention kernels to collocate prefill and decode phases, while LithOS [coppock2025lithos] atomizes kernels to pack execution fragments at the Texture Processing Cluster (TPC) level. While effective for utilization, these invasive techniques necessitate user-kernel modifications, risking behavioral inequivalence and complicating offline-to-online correctness guarantees. In contrast, DetShare avoids this semantic trap: by decoupling logical contexts from physical hardware, it dynamically remaps quotas at runtime to achieve elastic, fine-grained spatial multiplexing while strictly preserving execution determinism.
@@
The unique resource asymmetries of LLM inference have spurred specialized scheduling techniques. Recent work manages the prefill-decode imbalance via stall-free chunking (Sarathi-Serve [agrawal2024taming]) and fairness-aware phase balancing (FairBatching [lyu2025fairbatching]). Other efforts explore temporal scheduling through preemptible context migrations (Llumnix [sun2024llumnix]) and output-length-guided token-level preemption (TRAIL [shahoutdon]), or focus on QoS mitigation via adapter-aware caching (Chameleon [iliakopoulou2025chameleon]) and real-time latency attribution (LLMVisor [jinllmvisor], Niyama [goel2025niyama]). While these systems optimize application-level scheduling, DetShare provides a transparent, infrastructure-level foundation. Its extensible policy framework allows operators to plug in diverse algorithms while enforcing execution through strict, hardware-level isolation primitives.
@@
Frameworks like AvA [yu2020ava], rCUDA [duato2011enabling], and DCUDA [guo2019dcuda] leverage API remoting for disaggregated execution and flexible GPU kernel placement. Simultaneously, memory-centric systems like DeepUM [jung2023deepum], Zico [lim2021zico], and TGS [choi2022serving] optimize device utilization by reducing allocation overheads and shrinking active working-set footprints. These memory optimization and disaggregation techniques are entirely orthogonal and complementary to DetShare, presenting natural avenues for future integration to further maximize efficiency in shared environments.

### 17 Mar 2026 v2

@@ 3.3 Challenge-#3: Transparency and Practicality
However, many existing spatial sharing designs [15, 67] sacrifice this practicality for theoretical utilization gains. Approaches relying on static compiler analysis, operator fusion, or manual kernel rewriting introduce a high "Engineering Tax". By requiring deep introspection into model graphs or invasive source code modifications, these methods tightly couple the sharing substrate with specific model implementations. This coupling creates a brittle ecosystem where even minor updates to framework or model architecture can invalidate the optimization logic, necessitating expensive manual retuning.
@@
Crucially, this coupling imposes a burden beyond development—it fundamentally expands the verification surface. In hyperscale production environments, maintaining offline-to-online equivalence is non-negotiable. Mechanisms that modify user GPU kernels or the runtime environment effectively invalidate rigorous offline validation processes. This operational reality is corroborated by infrastructure practitioners at ByteDance [7], who note that "mechanisms requiring code rewriting are typically rejected not just for complexity, but because they mandate a prohibitive re-verification process for every framework or model update."
@@ 4 DetShare Design
This section details the design of DetShare, a system that reconciles execution determinism with high hardware utilization. Existing spatial sharing approaches often compromise semantic determinism by transparently altering block size to fit fragmented resources [9]. In contrast, DetShare adopts a transparent spatial multiplexing approach. The core design philosophy of DetShare is the strict decoupling of logical execution flows from physical resource provisioning. By maintaining kernel launch configurations as immutable invariants, DetShare ensures that resource elasticity never alters the arithmetic behavior of the application.
@@ 6.1 End-to-End Performance Evaluation
We compare DetShare against industry-standard hardware partitioning mechanisms (NVIDIA MPS and MIG), as well as state-of-the-art software-defined schedulers including Orion [67], Tick-Tock [71], and Salus [81]. Since the source code for LithOS [9] is not publicly available, we implemented a faithful reproduction of its scheduling logic based on its design principles to serve as a comparative baseline.
@@ 6.1.1 Colocated AI Model Training
High Compute Contention A ResNet-50 [33]
@@ B ResNet-101
C BERT-Large [13]
@@ D BERT-Large ResNet-50
Low Compute Contention 1 RNN [62]
2 DeepFM [29] ResNet-101
LLM Training E GLM-4-9B [23] (LoRA [36])
@@ 2 DeepFM [29] ResNet-101
CV + LLM Inference A ResNet Llama 3 [18]
B ResNet GPT-J [70]
@@ 6.2.1 Traces and Baselines
To demonstrate DetShare’s transparency and effectiveness at the infrastructure layer, we evaluate an unmodified instance of the state-of-the-art serving system, SGLang [85], running entirely on top of DetShare to serve the Llama-3.1-8B-Instruct model. We compare this against a baseline of SGLang executing on a default, exclusively provisioned GPU.
@@
  • Azure LLM Trace 2024 [66]: Highly stochastic request arrival times, testing general adaptability.
  • LongBench [3]: Long-context tasks imposing significant pressure on GPU memory capacity and bandwidth.
  • BurstGPT [72]: Extreme traffic spikes challenging the scheduler’s queue management.
@@ 7 Related Work
Temporal sharing remains the default mechanism for multi-tenant GPU execution. Prior systems have significantly enhanced this approach for DNN workloads by mitigating context-switching overheads through state reuse (Salus [81]), dynamically managing containerized or runtime memory footprints (Antman [77], TGS [75], Zico [48]), and employing predictive scheduling to bound inference tail latency (Clockwork [27], Gandiva [76]). However, these solutions inherently operate at the granularity of the entire GPU. Consequently, they enforce strict execution serialization and fail to overlap computation. This limitation prevents them from reclaiming idle capacity during memory-bound phases or bursty workloads, leaving significant hardware resources underutilized.
@@
Spatial sharing partitions hardware to enable concurrent execution. Foundational mechanisms include intra-context CUDA Streams [12] and NVIDIA’s static partitioning tools, MPS and MIG. Systems such as GSlice [15], GPUlets [8], and COLTI [53] build upon these primitives to construct higher-level spatio-temporal abstractions. Yet, because their underlying substrates rely on rigid, static partitioning, they cannot elastically adapt to the highly dynamic resource demands characteristic of modern workloads.
Recent literature pushes for even finer-grained multiplexing. POD-Attention [43] rewrites attention kernels to collocate prefill and decode phases, while LithOS [9] atomizes kernels to pack execution fragments at the Texture Processing Cluster (TPC) level. While effective for utilization, these invasive techniques necessitate user-kernel modifications, risking behavioral inequivalence and complicating offline-to-online correctness guarantees. In contrast, DetShare avoids this semantic trap: by decoupling logical contexts from physical hardware, it dynamically remaps quotas at runtime to achieve elastic, fine-grained spatial multiplexing while strictly preserving execution determinism.
@@
The unique resource asymmetries of LLM inference have spurred specialized scheduling techniques. Recent work manages the prefill-decode imbalance via stall-free chunking (Sarathi-Serve [1]) and fairness-aware phase balancing (FairBatching [52]). Other efforts explore temporal scheduling through preemptible context migrations (Llumnix [68]) and output-length-guided token-level preemption (TRAIL [63]), or focus on QoS mitigation via adapter-aware caching (Chameleon [37]) and real-time latency attribution (LLMVisor [41], Niyama [24]). While these systems optimize application-level scheduling, DetShare provides a transparent, infrastructure-level foundation. Its extensible policy framework allows operators to plug in diverse algorithms while enforcing execution through strict, hardware-level isolation primitives.
@@
Frameworks like AvA [80], rCUDA [17], and DCUDA [28] leverage API remoting for disaggregated execution and flexible GPU kernel placement. Simultaneously, memory-centric systems like DeepUM [42], Zico [48], and TGS [8] optimize device utilization by reducing allocation overheads and shrinking active working-set footprints. These memory optimization and disaggregation techniques are entirely orthogonal and complementary to DetShare, presenting natural avenues for future integration to further maximize efficiency in shared environments.